# MEt3R: Measuring Multi-View Consistency in Generated Images

### v1

@@ 5.3.1 Multi-View Generation
Following the validation of MEt3R in comparison to other metrics, we now benchmark our multi-view generation baselines on the test sequences (c.f. Sec. 5.1). In Tab. 1(a), we plot MEt3R against FID and KID along with the respective model size in terms of the number of parameters. We find that GenWarp achieves the worst consistency in terms of MEt3R, where the contents of the scene change drastically as we transition from one image to another, which can be qualitatively observed in Figs. 5,  1519. This behavior is expected since GenWarp generates one image at a time. Meanwhile, PhotoNVS performs slightly better than GenWarp but produces low-quality results, which are captured quantitatively by FID and KID. GenWarp and PhotoNVS cannot learn an expressive multi-view prior since they are single-view generation models, hindering their ability to produce 3D consistent results.
Conversely, diffusing multiple views at a time induces a stronger prior towards 3D consistency, as in MV-LDM, where we see an overall improvement in MEt3R. Among all evaluated methods, MV-LDM achieves the best trade-off between 3D consistency and novel view quality, both qualitatively and quantitatively. Moving further towards 3D consistency, DFM uses an underlying 3D representation and, therefore, produces consistent novel views by design, which is captured quantitatively in the form of better MEt3R scores than MV-LDM. However, this strong inductive bias comes with the drawback of blurry renderings far away from the ground-truth image distribution, as reflected by FID and KID. This highlights that MEt3R only focuses on 3D consistency irrespective of image content and can therefore complement standard image quality metrics well.
Figure 6: Feature similarity ablation. We compare MEt3R against versions of it that compare RGB projections via PSNR and SSIM. It can be seen that the PSNR versions give better scores to DFM than to real videos. We attribute this to their sensitivity to view-dependent effects, such as lighting. In contrast, MEt3R rates the real video best. Further, the standard deviation of PSNR and SSIM versions are much higher, also for real videos, indicating a lower signal-to-noise ratio.
MEt3R on multiple scales.
$\mathbf{256^{2}}$ $\mathbf{224^{2}}$ $\mathbf{128^{2}}$
GenWarp [31] 0.120 1.398 +$2.58\%$ +$74.70\%$$6.95\%$ +$80.27\%$
PhotoNVS [46] 0.069 0.479 +$3.86\%$ +$26.22\%$$1.78\%$ +$42.53\%$
MV-LDM (Ours) 0.036 0.405 +$3.88\%$$2.27\%$$16.63\%$ +$34.28\%$
DFM [36] 0.026 0.346 +$9.39\%$ +$52.73\%$$5.97\%$ +$70.15\%$
Real Video 0.022 0.181 +$2.44\%$ +$18.98\%$$47.80\%$ +$148.39\%$
@@ 5.3.2 Video Generation
A particular advantage of MEt3R is that it does not require camera poses to measure consistency, unlike TSED [46] and SED [46]. Thus, it is the first quantitative multi-view consistency metric that can be used on videos generated by video diffusion models.
Figure 7: MEt3R on generated videos. Per-image-pair plot for MEt3R across 48 frames and averaged across 100 sequences of RealEstate10K. For I2VGen-XL, we observe large inconsistencies initially as the inputs are out-of-distribution, followed by gradual improvement, indicating the inputs get closer to being in-distribution, also visible qualitatively in Figs. 1519. Meanwhile, Ruyi-Mini-7B shows several periodic spikes indicating abrupt inconsistencies throughout the video sequence, whereas MEt3R for SVD stays relatively low and smooth.
Table 1 shows the average MEt3R along with FID, KID, and FVD. Moreover, Fig. 7 shows the average MEt3R per image pair for I2VGen-XL [47], Ruyi-Mini-7B [35] and SVD[2] which clearly shows that SVD has better 3D consistency than Ruyi-Mini-7B and I2VGen-XL. However, SVD generates smoother and shorter camera trajectories, whereas Ruyi-Mini-7B and I2VGen-XL produce large motion at the expense of 3D consistency. Ruyi-Mini-7B is consistently worse, with periodic spikes, attributed to unstable camera motion and sudden 3D inconsistencies. For I2VGen-XL, as the inputs are out of distribution, MEt3R starts from a higher value followed by a gradual improvement as the model forces each progressing sample to be more in distribution while preserving similar global structures as in the initial input image. The resulting MEt3R scores correlate with visual judgment about the 3D consistency of the baseline video diffusion models, which can be observed in Figs. 1519 in the appendix.
@@ 5.4 Analyzing Alternative Similarities
Figure 8: Feature backbone ablation. We analyze the effect of different feature backbones on MEt3R. While DINOv2 [26] and MaskCLIP [48] can be employed as well, we found DINO features to lead to a more informative separation of models.
@@
[…]
@@
In Fig. 8, we evaluate MEt3R in combination with DINOv2 [26] and MaskCLIP [48] as alternatives to DINO [4] in the feature backbone. DINOv2 and MaskCLIP strongly compress the values in a tighter range, reducing the gap between extremely inconsistent and consistent generation. We find that DINO features provide a better separation of model performance and capture substantial inconsistencies more reliably, as seen from the random noise. Nevertheless, MEt3R is flexible with this design choice as better and more 3D consistent feature backbones can be used to improve the overall metric and to further reduce the lower bound.
@@ Acknowledgements
The supplementary materials are structured as follows. First, we provide detailed information about our multi-view latent diffusion model (MV-LDM) in Sec. A. Then, we provide details about the MEt3R metric in Sec. B. Finally, we present additional details on the multi-view generation baselines in Sec. C and their corresponding runtime statistics in Sec. D. Please also note our supplementary video, showcasing evaluations in motion.
@@ Appendix A Multi-View Latent Diffusion Model
Like CAT3D [10], our architecture is based on a multi-view 2D UNet shared across multiple input views with 3D self-attention at each UNet block. We initialize the UNet weights with Stable Diffuson 2.1 [28] and replace each attention layer with a 3D self-attention layer from MVDream [32] where each token from one view attends to all tokens from the other views. This accounts for 1.1B parameters for the multi-view UNet and 83.7M for the VAE. Due to memory and resource limitations, we fix the total number of concurrent views to 5, including the target and the conditioning views. Figure 9 shows the architecture of MV-LDM. We apply a VAE encoder and map the input images $(H\times W\times 3)$ into latent representation $(\frac{H}{8}\times\frac{W}{8}\times 4)$. For the low-resolution latent maps, we generate the ray encodings of shape $(\frac{H}{8}\times\frac{W}{8}\times 6)$, which consists of a 3-dimensional origin and a 3-dimensional direction vector in relative camera space and concatenate it along the channel dimension.
@@ A.2 Training and Evaluation with MEt3R
The training procedure follows DDPM [13], sampling a noise level $t$, applying that to all given latent images and training the network to predict the noise present in the image. We randomly select the conditioning views $N$ between 1 or 2 and the target views $M$ between 3 and 4, respectively, to allow for single and few-view novel view generation. We linearly vary the beta schedule from 0.0001 to 0.02 for the forward diffusion process and train MV-LDM for a total of 1.65M iterations with an effective batch size of 24 at resolution $256^{2}$. We use AdamW [22] optimizer with a constant learning rate of $2e^{-5}$. During sampling, the network can receive a combination of existing and pure noise images with camera ray encodings to perform conditional generation. The backward diffusion process is done with $\epsilon$-parameterization defined as the output of the MV-LDM $\boldsymbol{\epsilon}_{\theta}$ as,
@@
$\mathcal{L}=||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{z}_{t},\mathbf{c}_{t},t\right)||^{2}_{2}\textit{,}$ (6)
@@
We further test MEt3R with two sampling strategies, i.e., (1) autoregressively generating new target views and new anchors, conditioned on the previous anchor, and (2) using anchored sampling where we generate anchors first and then the rest as described in Sec. 4. Fig. 11 shows the average MEt3R plot per image-pair, showing the improvements with anchored sampling. For autoregressive sampling, we observe many diverging peaks that refer to several anchor-to-anchor transitions and accumulating errors. In contrast, for anchored sampling, the anchors are generated together first, followed by generating the rest. This limits error accumulation and results in fewer anchor-to-anchor transitions. Refer to Fig. 12 for a visual illustration of anchored and autoregressive sampling schemes.
@@
Appendix B Additional MEt3R Architectural Details
@@
Then, we extract the $x$, $y$, and $z$ coordinate maps from $\mathbf{X}_{canon}$ as $\mathbf{X},\mathbf{Y},\mathbf{Z}\in\mathbb{R}^{H\times W}$. Moreover, DUSt3R already implements this in their codebase, which we incorporate in MEt3R as shown in Alg. 1. The computed focal length $f_{x}$ and $f_{y}$, along with the principle point offsets $c_{x}$ and $c_{y}$, are used to form the projection matrix.
@@
2:$f_{x},f_{y}$
3:$\mathbf{Q}_{x}=\frac{\mathbf{U}\odot\mathbf{Z}}{\mathbf{X}}$triangleright $\odot$ is the Hadamard Product
4:$\mathbf{Q}_{y}=\frac{\mathbf{V}\odot\mathbf{Z}}{\mathbf{Y}}$
5:$f_{x}=median(\mathbf{Q}_{x})$triangleright Across spatial dimension
6:$f_{y}=median(\mathbf{Q}_{y})$
@@
Figure 13: Fixed vs. adjusted projection matrix. With fixed focal length, the projection area varies across different scales of DUSt3R point maps. We automatically adjust the focal length for each example pair to allow maximal projection and, therefore, more pixels for evaluating feature similarity.
@@
Appendix C Additional Details on Multi-View Generation Models
@@ Appendix D Runtime
Table 3: Runtime comparison. We report the runtime in seconds for all the baselines for generating a full video sequence comprising 80 frames. MV-LDM and GenWarp achieve the fastest sampling, followed by DFM and then PhotoNVS.

### v2

@@ 5.3.1 Multi-View Generation
Following the validation of MEt3R in comparison to other metrics, we now benchmark our multi-view generation baselines on the test sequences (c.f. Sec. 5.1). In Tab. 1(a), we plot MEt3R against FID [13] and FVD [40] along with the respective model size in terms of the number of parameters. We find that GenWarp [32] achieves the worst consistency in terms of MEt3R, where the contents of the scene change drastically as we transition from one image to another, which can be qualitatively observed in Figs. 5,  1721. This behavior is expected since GenWarp generates one image at a time. Meanwhile, PhotoNVS [48] performs slightly better than GenWarp but produces low-quality results, which the FID captures quantitatively. GenWarp and PhotoNVS cannot learn an expressive multi-view prior since they have a single-sized context window, hindering their ability to produce consistent 3D results.
Conversely, diffusing multiple views at a time induces a stronger prior towards 3D consistency, as in MV-LDM, where we see an overall improvement in MEt3R. Among all evaluated methods, MV-LDM achieves the best trade-off between 3D consistency and novel view quality, both qualitatively and quantitatively. Moving further towards 3D consistency, DFM [38] uses an underlying 3D representation and produces consistent novel views by design, which is captured quantitatively in the form of better MEt3R scores than MV-LDM. However, this strong inductive bias comes at the cost of blurry renderings pushing further from the ground-truth distribution, as reflected by the FID, and achieves a higher signal-to-noise ratio, as indicated by FWS (PSNR). This highlights that MEt3R only focuses on 3D consistency irrespective of image content and can, therefore, complement standard image quality metrics well.
@@ 5.3.2 Video Generation
A particular advantage of MEt3R is that it does not require camera poses, unlike TSED [48] and SED [48], where it can be used directly on generated videos to measure consistency similar to FWS (PSNR).
Figure 6: Average pairwise MEt3R on generated videos. Per-image-pair plot for MEt3R across 48 frames and averaged across 100 sequences of RealEstate10K [51]. We find that SVD [2] achieves the best MEt3R score, followed by Ruyi-Mini-7B [36] and I2VGenXL [49].
Table 1 shows the average MEt3R, FID, FWS (PSNR), and FVD. Moreover, Fig. 6 shows the average MEt3R per image pair for I2VGen-XL [49], Ruyi-Mini-7B [36] and SVD[2] which shows that SVD has better 3D consistency than Ruyi-Mini-7B and I2VGen-XL. However, SVD generates smoother and shorter camera trajectories, whereas Ruyi-Mini-7B and I2VGen-XL produce large motion at the expense of 3D consistency. For I2VGen-XL, as the inputs are out of distribution, MEt3R starts from a higher value followed by a gradual improvement as the model forces each progressing sample to be more in distribution while preserving similar global structures as in the initial input image. This behavior is also qualitatively visible in Figs. 1721. Meanwhile, Ruyi-Mini-7B shows several spikes indicating abrupt inconsistencies throughout the video sequence, which are attributed to unstable camera motion. Furthermore, in Tab. 1, FWS (PSNR) gives the lowest score to Ruyi-Mini-7B, indicating a slightly noisy generation even though it is more 3D consistent than I2VGen-XL.
Figure 7: Object-level evaluation on GSO [9]. Average pairwise MEt3R on 30 examples, each consisting of $360^{\circ}$ rotation around the object with loop closure, i.e., at $360^{\circ}$, we evaluate the first and the last frame. We find that SyncDreamer [22] achieves the best in MEt3R, followed by VideoMV [53] and EpiDiff [16] while respecting the lower bound.
5.3.3 Object-Level Generation
Lastly, we evaluate object-level diffusion models on GSO [9] dataset. We consider EpiDiff [16], VideoMV [53] and SyncDreamer [22] as the baselines and are evaluated on $360^{\circ}$ camera rotation spread across 16 frames on 30 examples. We find that MEt3R can differentiate models with various levels of inconsistencies, which we report in Fig. 7. Although the baselines show good visual quality, consistency varies heavily. Notice how MEt3R captures the slightly increasing inconsistency for EpiDiff when moving further away from the condition, which suggests that the strength of conditioning plays an important role in producing better consistency (c.f. Fig. 4 and Sec. 5.2). Further qualitative results are provided in Figs. 2225 in the appendix.
@@ 5.4 Analyzing Alternative Similarities
Figure 8: Feature similarity ablation. We compare MEt3R against versions of it that compare RGB projections via PSNR and SSIM. It can be seen that unlike MEt3R, PSNR and SSIM give better scores to DFM than to real videos. We attribute this to their sensitivity to view-dependent effects, such as lighting. Note that for real videos, the standard deviation of PSNR and SSIM are much higher, indicating a lower signal-to-noise ratio.
@@
Instead of projecting features onto a shared view, staying in RGB space would enable the use of classical image quality metrics such as PSNR and SSIM. Fig. 8 provides a comparison of such variants $\mathrm{MEt3R_{PSNR}}$ and $\mathrm{MEt3R_{PSNR}}$ with MEt3R. While a reasonable negative correlation can be observed, DFM [38] outperforms the ground-truth video w.r.t. these metrics. We attribute this to the bias of PSNR and SSIM to blur, which is apparent in novel views generated by DFM due to its low resolution and reliance on pixelNeRF [47] acting as an architectural bottleneck. In contrast, real videos exhibit view-dependent effects, including brightness variations and reflections, to which PSNR and SSIM are highly sensitive. With MEt3R, we aim to abstract from these pixel-level inconsistencies and instead provide a metric that robustly measures the 3D consistency of generative approaches. Therefore, we opt for similarities in a suitable feature space.
@@
In Fig. 9, we evaluate MEt3R in combination with DINOv2 [27] and MaskCLIP [50] as alternatives to DINO [4] in the feature backbone. DINOv2 and MaskCLIP strongly compress the values in a tighter range, reducing the gap between extremely inconsistent and consistent generation. We find that DINO features provide a better separation of model performance and capture substantial inconsistencies more reliably, as seen from the random noise. Nevertheless, MEt3R is flexible with this design choice as better and more 3D consistent feature backbones can improve and reduce the lower bound further.
Figure 9: Feature backbone ablation. We analyze the effect of different feature backbones on MEt3R. While DINOv2 [27] and MaskCLIP [50] can be employed as well, we found DINO [4] features to lead to a more informative separation of models.
@@ Acknowledgements
The supplementary materials are structured as follows. First, we provide detailed information about our multi-view latent diffusion model (MV-LDM) in Sec. A. Then, we compare MEt3R with the FWS variants in Sec. B, followed by a discussion on MEt3R metric in Sec. C. Finally, we present additional details on the multi-view generation baselines in Sec. D and their corresponding runtime statistics in Sec. E. Please also note our supplementary video, showcasing evaluations in motion.
@@ Appendix A Multi-View Latent Diffusion Model
Like CAT3D [11], our architecture is based on a multi-view 2D UNet shared across multiple input views with 3D self-attention at each UNet block. We initialize the UNet weights with Stable Diffusion 2.1 [28] and replace each attention layer with a 3D self-attention layer from MVDream [33] where each token from one view attends to all tokens from the other views. This accounts for 1.1B parameters for the multi-view UNet and 83.7M for the VAE. Due to memory and resource limitations, we fix the total number of concurrent views to 5, including the target and the conditioning views. Figure 10 shows the architecture of MV-LDM. We apply a VAE encoder and map the input images $(H\times W\times 3)$ into latent representation $(\frac{H}{8}\times\frac{W}{8}\times 4)$. For the low-resolution latent maps, we generate the ray encodings of shape $(\frac{H}{8}\times\frac{W}{8}\times 6)$, which consists of a 3-dimensional origin and a 3-dimensional direction vector in relative camera space and concatenate it along the channel dimension.
@@ A.2 Training and Evaluation with MEt3R
The training procedure follows DDPM [14], sampling a noise level $t$, applying that to all given latent images and training the network to predict the noise present in the image. We randomly select the conditioning views $N$ between 1 or 2 and the target views $M$ between 3 and 4, respectively, to allow for single and few-view novel view generation. We linearly vary the beta schedule from 0.0001 to 0.02 for the forward diffusion process and train MV-LDM for a total of 1.65M iterations with an effective batch size of 24 at resolution $256^{2}$. We use AdamW [23] optimizer with a constant learning rate of $2e^{-5}$. During sampling, the network can receive a combination of existing and pure noise images with camera ray encodings to perform conditional generation. The backward diffusion process is done with $\epsilon$-parameterization defined as the output of the model $\boldsymbol{\epsilon}_{\theta}$:
@@
[…]
@@
We further test MEt3R with two sampling strategies, i.e., (1) autoregressively generating new target views and new anchors, conditioned on the previous anchor, and (2) using anchored sampling where we generate anchors first and then the rest as described in Sec. 4. Fig. 12 shows the average MEt3R plot per image-pair, showing the improvements with anchored sampling. We observe many diverging peaks attributed to several anchor-to-anchor transitions and accumulating errors for autoregressive sampling. For anchored sampling, the anchors are generated together first, followed by generating the rest. This limits the error accumulation and allows for fewer anchor-to-anchor transitions. We refer to Fig. 13 for a visual illustration of anchored and autoregressive sampling schemes.
@@
MEt3R on multiple scales.
$\mathbf{256^{2}}$ $\mathbf{224^{2}}$ $\mathbf{128^{2}}$
[…]
Appendix B Comparison of FWS Variants
Other metrics based on flow warping score (FWS) have been introduced to measure consistency, which uses optical flow, e.g., RAFT [37]. Given a pair of images, it first computes optical flow, which is used to warp one image into the other. Then, metrics such as SSIM, LPIPS, PSNR, and RMSE are computed to quantify multi-view consistency.
In Tab. 3, we evaluate both multi-view and video generation baselines on FWS and MEt3R. We find that most variants, including PSNR, SSIM, and RMSE, rank DFM better than real video among multi-view generation methods because of their sensitivity to blur, pixel-level perturbation, and noise (c.f. Sec. 5.3 and 5.4). Meanwhile, MEt3R and FWS (LPIPS) ignore such perturbations and rely on feature and perceptual similarity, respectively. However, as shown in Fig. 14, FWS (LPIPS) suffers at higher frame distances between input pairs, where DFM, GenWarp, and MV-LDM can score better than real videos.
FWS
MEt3R $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE $\downarrow$
GenWarp  [32] 0.120 21.41 0.716 0.200 0.097
PhotoNVS [48] 0.069 25.10 0.779 0.137 0.060
MV-LDM (Ours) 0.036 28.46 0.851 0.095 0.044
DFM [38] 0.026 39.56 0.948 0.082 0.011
Real Video 0.022 33.52 0.924 0.075 0.026
I2VGen-XL [49] 0.050 28.62 0.844 0.107 0.044
Ruyi-Mini-7B [36] 0.047 28.01 0.831 0.133 0.043
SVD [2] 0.032 29.93 0.890 0.079 0.038
Real Video 0.022 33.60 0.925 0.074 0.026
Table 3: Comparison of flow warping scores (FWS) with MEt3R. We report the results on multi-view and video generation methods, with FWS variants including PSNR, SSIM, LPIPS, and RMSE.
Figure 14: MEt3R comparison to existing Flow Warping Score (FWS). For a gap (frame distance) of 10 between the image pair, MET3R is more robust and does not violate the lower bound, unlike FWS.
Appendix C Additional MEt3R Architectural Details
@@
Then, we extract the $x$, $y$, and $z$ coordinate maps from $\mathbf{X}_{canon}$ as $\mathbf{X},\mathbf{Y},\mathbf{Z}\in\mathbb{R}^{H\times W}$. Moreover, DUSt3R already implements this in their codebase, which we incorporate in MEt3R as shown in Alg. 1. The computed focal length $f_{x}$ and $f_{y}$, along with the principal point offsets $c_{x}$ and $c_{y}$, are used to form the projection matrix.
@@
[…]
@@
Figure 15: Fixed vs. adjusted projection matrix. With fixed focal length, the projection area varies across different scales of DUSt3R [42] point maps. We automatically adjust the focal length for each example pair to allow maximal projection and, therefore, more pixels for evaluating feature similarity.
@@
Appendix D Additional Details on Multi-View Generation Models
@@ Appendix E Runtime
Table 4: Runtime comparison. We report the runtime in seconds for all the baselines for generating a full video sequence comprising 80 frames. MV-LDM and GenWarp [32] achieve the fastest sampling, followed by DFM [38] and then PhotoNVS [48].
@@
Figure 22: Qualitative results on GSO. A $360^{\circ}$ rendering of the GSO object Elephant.
Figure 23: Qualitative results on GSO. A $360^{\circ}$ rendering of the GSO object Alarm.
Figure 24: Qualitative results on GSO. A $360^{\circ}$ rendering of the GSO object Blocks.
Figure 25: Qualitative results on GSO. A $360^{\circ}$ rendering of the GSO object Cream.